OpenStudy (anonymous):

Which is the inverse of y = 1/2 x + 3? A y = 2 x − 3 B y = 2 x − 6 C y = –3x + 2 D y = – 1/ 2 x − 3

OpenStudy (hoblos):

y = 1/2 x + 3 to find the inverse replace x by y and y by x x=1/2y +3 1/2y = x-3 y = 2x-6